# Order-Optimal Instance-Dependent Bounds for Offline Reinforcement Learning with Preference Feedback

### v1

@@ 5 Extension to $(\varepsilon,\delta)$-Differential Privacy (DP)
In this section, we extend our algorithm RL-LOW to be amenable to $(\varepsilon,\delta)$-differential privacy. To formalize our results, we provide the definition of $(\varepsilon,\delta)$-DP, following Dwork et al., (2014). We say that two sets of preference labels, $\sigma:=\{\sigma_{i}\}_{i=1}^{n}$ and $\sigma^{\prime}:=\{\sigma^{\prime}_{i}\}_{i=1}^{n}$ are neighboring if there exists $s\in[n]$ such that $\sigma_{s}\neq\sigma^{\prime}_{s}$ and $\sigma_{j}=\sigma^{\prime}_{j}$ for all $j\neq s$.
@@
(Differential Privacy) Fix any label-free dataset $\{(s_{i},a^{(1)}_{i},a^{(2)}_{i})\}_{i=1}^{n}$. A (randomized) policy :$\mathcal{M}:\{0,1\}^{n}\rightarrow\mathcal{A}^{S}$ (that takes as inputs a set of labels and outputs a set of actions, one for each state) satisfies $(\varepsilon,\delta)$-DP if for all neighboring labels $\sigma:=\{\sigma_{i}\}_{i=1}^{n}$ and $\sigma^{\prime}:=\{\sigma^{\prime}_{i}\}_{i=1}^{n}$,
$\mathbb{P}(\mathcal{M}(\sigma)\in\mathcal{Z})\leq e^{\varepsilon}\,\mathbb{P}(\mathcal{M}(\sigma^{\prime})\in\mathcal{Z})+\delta\qquad\forall\,\mathcal{Z}\subset\mathcal{A}^{S}.$ (18)
Note that Definition 5.1 primarily concerns protecting the privacy of users’ labels. In particular, the DP mechanism guarantees that any alteration in a user’s label must not substantially affect the output of our algorithm. Otherwise, there exists a risk that a user’s label might be inferred through the algorithm’s output. Our definition of differential privacy (DP) aligns with that of Chowdhury et al., (2024).
We now adapt our RL-LOW to $(\varepsilon,\delta)$-DP by using the Gaussian mechanism  (Dwork et al.,, 2014). Firstly, we introduce the private version of the empirical success rate (analogous to $B_{k,i,j}$ in (5)) which we denote as
$\tilde{B}_{k,i,j}\coloneqq\frac{1}{nN_{k,i,j}}\sum_{\iota=1}^{n}\sigma_{\iota}\mathbbm{1}\{s_{\iota}=k,a_{\iota}^{(1)}=i,a_{\iota}^{(2)}=j\}+\tilde{\xi}_{k,i,j}\ \ \forall\,k\in\mathcal{S},(i,j)\in\mathcal{A}^{2}\text{ with }{N_{k,i,j}>0},$
where $\tilde{\xi}_{k,i,j}$ is an independent (across $k$, $i$, and $j$) Gaussian noise with zero mean and variance $\frac{2\log(1.25/\delta)}{(\varepsilon nN_{k,i,j})^{2}}$, and we let $\tilde{B}_{k,j,i}\coloneqq 1-\tilde{B}_{k,i,j}$. If $N_{k,i,j}=N_{k,j,i}=0$, we define $\tilde{B}_{k,i,j}=\tilde{B}_{k,j,i}=0$. Again, analogously to the operation in (6), we clip $\tilde{B}_{k,i,j}$ to form
@@
where $w^{(k,i,j)}$ is defined in Definition 3.1. Finally, the empirical best action is $\hat{i}_{k}\in\tilde{\mathcal{I}}_{k}:=\{i\in\mathcal{A}:\tilde{r}_{k,i,j}\geq 0,\forall\,j\neq i\}$. A similar argument as Proposition 3.2 shows that $\hat{i}_{k}$ exists; see details in Appendix C for the details. The algorithm described above is a differentialy private version of the RL-LOW policy and hence, it is named DP-RL-LOW.
@@
Given privacy parameters $\varepsilon,\delta>0$, DP-RL-LOW satisfies $(\varepsilon,\delta)$-DP.
The proof of Proposition 5.2 follows exactly along the lines of the proof of Dwork et al., (2014, Theorem A.1) and is omitted. We next upper bound the expected simple regret of DP-RL-LOW.
@@
$\!\!H_{\rm DP}^{(\varepsilon,\delta)}(v)\!=\!\max_{k\in\mathcal{S},i\in\mathcal{A}:i\neq i^{*}_{k}}\frac{\sqrt{\log(\frac{1.25}{\delta})\gamma^{\rm DP}_{k,i}}}{\sqrt{\varepsilon}\Delta_{k,i}}\;\;\text{and}\;\;\gamma_{k,i}^{\rm DP}\!=\!{\sum_{k^{\prime},i^{\prime},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\left(\frac{w^{(k,i,i^{*}_{k})}_{k^{\prime},i^{\prime},j^{\prime}}}{N_{k^{\prime},i^{\prime},j^{\prime}}}\right)^{2}}.$ (22)
@@
The limiting statement in (23) implies that DP-RL-LOW has the same order of the exponential decay rate as its non-differentially privacy counterpart RL-LOW when $n$ is sufficiently large; in particular, $n>{(H_{\rm DP}^{(\varepsilon,\delta)}(v))^{2}}/H(v)$ suffices to nullify the effect of the privacy requirement. In other words, in the sense of the exponent, privacy comes “for free” for sufficiently large offline datasets.
6 Numerical Simulations
Figure 1: Comparison of RL-LOW and DP-RL-LOW to Pessimistic MLE (Zhu et al.,, 2023) on average simple regret and standard deviation (shaded area). In the left figure, we set $\delta=0.2$ and $\varepsilon=2.0$ for DP-RL-LOW. In the right figure, we set $n=300$ for all policies, and $\varepsilon$ does not apply to RL-LOW.
@@ 6 Numerical Simulations
In this section, we present some numerical simulations of our algorithm RL-LOW and and its differentially private counterpart DP-RL-LOW. We compare them to the state-of-the-art (non-private algorithm) Pessimistic MLE developed by Zhu et al., (2023). We conduct the experiments on a synthetic dataset. Specifically, we set the number of states $S=2$, the number of actions $A=10$, the dimensionality of the data $d=5$, the unknown parameter vector $\theta=[1,1,1,1,1]^{\top}$, and the state distribution $\rho=[0.4,0.6]$. The feature vector of each action is generated as follows: For the $i$-th action of state $k\in\{1,2\}$, we first uniformly generate a $d$-dimentional vector $\phi^{\prime}(k,i)$ with all non-negative elements and $\lVert\phi^{\prime}(k,i)\rVert_{1}=1$. Then, for each state $k\in\{1,2\}$, we set the feature vector of $i$-th action as $\phi(k,i)=\phi^{\prime}(k,i)-0.01(i-1)\theta$. That is, in both state $1$ and $2$, the best action is the first action, and the suboptimality gap of the $i$-th action is $0.05i$. In addition, for both states $k\in\{1,2\}$ and $i<j$, we set $N_{k,i,j}=\frac{1}{A(A-1)}$, i.e., the proportions of comparisons for this instance are uniform. In the simulation, we use $\lceil N_{k,i,j}n\rceil$ as the number of samples involved in the comparison between actions $i$ and $j$ under state $k$.
@@
As for the hyperparameters of Pessimistic MLE, we follow the default setting of Zhu et al., (2023, Section 3). In addition, Pessimistic MLE only works under the assumption that $\langle 1,\theta\rangle=0$ (Zhu et al.,, 2023, Assumption 2.1). Therefore, when running the experiments of Pessimistic MLE, we further set $d=6$, $\theta=[1,1,1,1,1,-5]$ and the $6$-th element of each feature vector is set to 00. Then, this new instance is mathematically equivalent to the original instance and additionally satisfies (Zhu et al.,, 2023, Assumption 2.1) which is needed for Pessimistic MLE.
@@
Lastly, as shown in Figure 1 (right), it is evident that achieving comparable performance between RL-LOW and DP-RL-LOW may necessitate substantially larger sample sizes $n$ when considering small privacy parameters of $\varepsilon$ and $\delta$. This observation is again consistent with our theoretical findings in Theorem 5.3.
7 Concluding Remarks
This paper addresses the problem of offline RL with preference feedback, aiming to determine the optimal action for each state to minimize the simple regret. We introduced a novel algorithm, RL-LOW, which achieves a simple regret of $\exp(-\Omega(n/H(v))$, where $n$ represents the number of data samples and $H(v)$ characterizes an instance-dependent hardness parameter related to the suboptimality gaps of each action. Additionally, we established a first-of-its-kind instance-dependent lower bound for offline RL with preference feedback, demonstrating the order-wise optimality of RL-LOW (in the exponential decay rate) through the matching of lower and upper bounds on the simple regret. To address privacy concerns, we extended RL-LOW to be amenable to $(\varepsilon,\delta)$-differential privacy, revealing that the hardness parameter $H(v)$ remains unchanged in the asymptotic regime as $n$ tends to infinity. This underscores RL-LOW’s effectiveness and robustness in preserving the privacy of observed rewards. Our focus on establishing instance-dependent bounds sets this work apart from previous research that focuses primarily on worst-case regret analyses in offline RL with preference feedback.
Some interesting directions for future research include extending our work to incorporate general reward functions (Zhan et al.,, 2024). In particular, a natural question concerns whether or not there exist an algorithm that is instance-dependent and order-optimal for general reward functions?
Supplementary Material for
“Order-Optimal Instance-Dependent Bounds for Offline Reinforcement Learning with Preference Feedback”
Appendix A Useful facts
@@
Definition A.1 (SubGaussian norm).
@@
Definition A.2 (Variance proxy).
@@
Lemma A.3 (Linear combination of subGaussian random variables).
Let $X_{1},\ldots,X_{n}$ be independent subGaussian random variables, where the variance proxy of $X_{i}$ is $\sigma_{i}^{2}$ Then, for any $a_{1},\dots,a_{n}\in\mathbb{R}$, the random variable $\sum_{i=1}^{n}a_{i}X_{i}$ is a subGaussian random variable with variance proxy $\sigma^{2}=\sum_{i=1}^{n}a_{i}^{2}\sigma_{i}^{2}$.
Lemma A.4.
@@
Lemma A.5.
(Adapted from (Vershynin,, 2018, Proposition 2.5.2)) For any subGaussian random variable $X$,
@@
Appendix B Proof of Proposition 2.3
Lemma B.1.
@@
which is a contradiction to that $v$ is inconsistent. This completes the proof of Lemma B.1. ∎
@@
Fix any inconsistent instance $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$. By Lemma B.1, there exists $(\underline{k},\underline{i})\in\mathcal{S}\times\mathcal{A}$ with $\underline{i}\neq i^{*}_{\underline{k}}$ such that
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{v}^{\pi}\left[\sum_{k\in\mathcal{S}}\rho_{k}\left(\max_{j\in\mathcal{A}}\langle\phi(k,j)-\phi(k,\hat{i}_{k}),\theta\rangle+\max_{j\in\mathcal{A}}\langle\phi(k,j)-\phi(k,\hat{i}_{k}),\theta^{\prime}\rangle\right)\right]$ (26)
@@
Appendix C The existence of $\hat{i}_{k}$ in RL-LOW and DP-RL-LOW
@@
Lemma C.1.
Fix any consistent instance $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$. Recall the definitions of $w^{(k,i,j)}$ and $\gamma_{k,i,i^{*}_{k}}$ in Eqn. (7) and Eqn. (9), respectively. Then, for any $(k,i)\in\mathcal{S}\times\mathcal{A}$ with $i\neq i^{*}_{k}$,
@@
$[\phi(k,i)-\phi(k,j)]_{\mathcal{G}}=\sum_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}u_{k^{\prime},i^{\prime},j^{\prime}}[\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime})]_{\mathcal{G}}.$
@@
By Lemma C.1, we get that under RL-LOW, for any $(k,i,j,j_{2})\in\mathcal{S}\times\mathcal{A}^{3}$,
@@
which implies that |^$\lvert\hat{\mathcal{I}}_{k}\rvert\geq 1$ under RL-LOW. This completes the proof of Proposition 3.2. ∎

### v2

@@ 6 Extension to $(\varepsilon,\delta)$-Differential Privacy (DP)
In this section, we extend our algorithm RL-LOW to be amenable to $(\varepsilon,\delta)$-differential privacy with labels, and we provide a motivational example of this extension in Appendix A. To formalize our results, we provide the definition of $(\varepsilon,\delta)$-DP, following Dwork et al. (2014). We say that two sets of preference labels, $\sigma:=\{\sigma_{i}\}_{i=1}^{n}$ and $\sigma^{\prime}:=\{\sigma^{\prime}_{i}\}_{i=1}^{n}$ are neighboring if there exists $s\in[n]$ such that $\sigma_{s}\neq\sigma^{\prime}_{s}$ and $\sigma_{j}=\sigma^{\prime}_{j}$ for all $j\neq s$.
@@
(Differential Privacy with labels) Fix any label-free dataset $\{(s_{i},a^{(1)}_{i},a^{(2)}_{i})\}_{i=1}^{n}$. A (randomized) algorithm $\mathcal{M}:\{0,1\}^{n}\rightarrow\mathcal{A}^{S}$ (that takes as inputs a set of labels and outputs a set of actions, one for each state) satisfies $(\varepsilon,\delta)$-DP if for all neighboring labels $\sigma:=\{\sigma_{i}\}_{i=1}^{n}$ and $\sigma^{\prime}:=\{\sigma^{\prime}_{i}\}_{i=1}^{n}$ it holds $\forall\,\mathcal{Z}\subset\mathcal{A}^{S}$
$\mathbb{P}(\mathcal{M}(\sigma)\in\mathcal{Z})\leq e^{\varepsilon}\,\mathbb{P}(\mathcal{M}(\sigma^{\prime})\in\mathcal{Z})+\delta.$ (18)
Note that Definition 6.1 primarily concerns protecting the privacy of users’ labels. In particular, the DP mechanism guarantees that any alteration in a user’s label must not substantially affect the output of our algorithm. Otherwise, there exists a risk that a user’s label might be inferred through the algorithm’s output. Our definition of differential privacy (DP) aligns with that of Chowdhury et al. (2024).
We now adapt our RL-LOW to $(\varepsilon,\delta)$-DP by using the Gaussian mechanism  (Dwork et al. 2014). Firstly, we introduce the private version of the empirical success rate (analogous to $B_{k,i,j}$ in Eqn. (5)), i.e., for all $k\in\mathcal{S}$ and $i,j\in\mathcal{A}$ and $N_{k,i,j}>0$,
$\tilde{B}_{k,i,j}\coloneqq\frac{1}{nN_{k,i,j}}\sum_{\iota=1}^{n}\sigma_{\iota}\mathds{1}\{s_{\iota}=k,a_{\iota}^{(1)}=i,a_{\iota}^{(2)}=j\}+\tilde{\xi}_{k,i,j},$
where $\tilde{\xi}_{k,i,j}$ is an independent (across $k$, $i$, and $j$) Gaussian noise with zero mean and variance $\frac{2\log(1.25/\delta)}{(\varepsilon nN_{k,i,j})^{2}}$, and we let $\tilde{B}_{k,j,i}\coloneqq 1-\tilde{B}_{k,i,j}$. If $N_{k,i,j}=N_{k,j,i}=0$, we define $\tilde{B}_{k,i,j}=\tilde{B}_{k,j,i}=0$. Again, analogously to the operation in Eqn. (6), we clip $\tilde{B}_{k,i,j}$ to form
@@
where $w^{(k,i,j)}$ is defined in Definition 3.1. Finally, the empirical best action is $\hat{i}_{k}\in\tilde{\mathcal{I}}_{k}:=\{i\in\mathcal{A}:\tilde{r}_{k,i,j}\geq 0,\forall\,j\neq i\}$. A similar argument as Proposition 3.2 shows that $\hat{i}_{k}$ exists; see details in Appendix F for the details. The algorithm described above is a differentially private version of RL-LOW and hence, it is named DP-RL-LOW.
@@
Given privacy parameters $\varepsilon,\delta\in(0,1)$, DP-RL-LOW satisfies $(\varepsilon,\delta)$-DP.
The proof of Proposition 6.2 follows exactly along the lines of the proof of Dwork et al. (2014, Theorem A.1) and is omitted. Then, the upper bound is as follows.
@@
$H_{\rm DP}^{(\varepsilon,\delta)}(v)\!=\!\max_{k\in\mathcal{S},i\in\mathcal{A}:i\neq i^{*}_{k}}\frac{\sqrt{\log(\frac{1.25}{\delta})\gamma^{\rm DP}_{k,i}}}{\sqrt{\varepsilon}\Delta_{k,i}},\;\;\text{and}\;$ (22)
$\gamma_{k,i}^{\rm DP}\!=\!{\sum_{k^{\prime},i^{\prime},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\left(\frac{w^{(k,i,i^{*}_{k})}_{k^{\prime},i^{\prime},j^{\prime}}}{N_{k^{\prime},i^{\prime},j^{\prime}}}\right)^{2}}.$ (23)
@@
The limiting statement in Eqn. (24) implies that DP-RL-LOW has the same order of the exponential decay rate as its non-differentially privacy counterpart RL-LOW when $n$ is sufficiently large; in particular, $n>{(H_{\rm DP}^{(\varepsilon,\delta)}(v))^{2}}/H(v)$ suffices to nullify the effect of the privacy requirement. In other words, in the sense of the exponent, privacy comes “for free” for sufficiently large offline datasets.
In addition, we derive the worst-case upper bound of DP-RL-LOW in Appendix I, which yields the form of $O(\frac{1}{\sqrt{n}}+\frac{\sqrt{\log(1/\delta)}}{\varepsilon n})$ for the dependency in $n$, $\varepsilon$ and $\delta$. This again implies that privacy comes for free for sufficiently large offline dataset.
7 Related Works Beyond PbRL
Offline RL without Human Feedback The domain of offline RL has been extensively researched over an extended period. Here, We focus on the recent works. Chen and Jiang (2019) revisits and provides theoretical insights into the essential but underexplored assumptions of mild distribution shift and strong representation conditions in value-function approximations, advancing their necessity and applicability.  Xie et al. (2021) bridges the gap between online and offline reinforcement learning by introducing the policy finetuning problem, proposing algorithms that leverage a reference policy close to the optimal policy to achieve sample-efficient learning in episodic MDPs.  Yin et al. (2022) investigates the statistical limits of offline reinforcement learning using linear models, introducing the variance-aware pessimistic value iteration method to improve learning bounds with offline data. More recently, Wang et al. (2022) enhances the understanding of gap-dependent sample complexity in offline reinforcement learning, demonstrating improved rates under specific policy coverage conditions and providing algorithms nearly matching lower bounds. Similarly,  Nguyen-Tang et al. (2023) investigated gap-dependent analysis for offline RL, providing both gap-dependent upper and lower bounds for performance with linear function approximation.
Overall, our study identifies a significant oversight in previous research: the absence of human feedback consideration. Consequently, our work represents the inaugural investigation into instance-dependent bounds within the context of offline reinforcement learning incorporating human feedback with pairwise comparisons.
Dueling Bandits The Dueling Bandits problem was first introduced by Yue and Joachims (2009), sparking a substantial body of subsequent research on the topic. In this section, we highlight some relevant works. Inspired by the classical contextual bandits problem, Dudík et al. (2015) extend the framework of duel bandits into a contextual setting, and they propose a new concept of von Neumann winner, a game-theoretic solution concept that addresses limitations of the Condorcet winner, along with three efficient algorithms for its online learning and approximation from data. In contrast,  Saha (2021) explore a distinct aspect of contextual dueling bandits through their proposed Subsetwise-Preference Feedback Model, and the author presents two algorithms for pairwise preferences, achieving near-optimal regret bounds, and extending the analysis to general subsetwise preferences, demonstrating that the fundamental performance limits remain consistent regardless of the subset size. However, this study mainly focuses on the worst-case analysis. More recently,  Di et al. (2024) addressed the contextual dueling bandits with adversarial feedback, proposing a robust algorithm using uncertainty-weighted maximum likelihood estimation. Nonetheless, this work focuses on the adversarial setting, whereas our work examines the stochastic setting.
8 Concluding Remarks
This paper studies offline RLHF with pairwise comparisons, aiming to minimize simple regret by identifying the optimal action per state. We propose new algorithms achieving simple regret of the form $\exp(-\Omega(n/H(v)))$, where $n$ is the sample size and $H(v)$ captures instance-dependent hardness from suboptimality gaps. We also establish the first instance-dependent lower bound for this setting, matching our upper bound and proving exponential-rate optimality. To ensure privacy, we adapt our method to be $(\varepsilon,\delta)$-differentially private, showing that $H(v)$ remains asymptotically unchanged as $n\to\infty$. By establishing instance-dependent bounds of exponential convergence, our results close a gap in prior works that focused on worst-case regret.
Acknowledgements
This research work was funded by two Singapore Ministry of Education Academic Research Fund Tier 1 grants (A-8000980-00-00 and A-8002934-00-00).
Appendix A Details on Motivation
A.1 A Motivational Example of Label-DP
In the development of question-answering (QA) systems, a common approach for improving response quality involves engaging users in a labeling process where they are asked to provide preference labels. Specifically, users evaluate pairs of system-generated responses to a given question and indicate which response they prefer. This method, often referred to as pairwise preference labeling, is integral to training RLHF algorithms that aim to optimize the relevance and utility of answers provided by QA systems.
Given our understanding of the nature of this process, our research emphasizes the importance of protecting the confidentiality of user-submitted preference labels. Without any concerted attempt to protect privacy, these labels, which directly reflect individual opinions or biases toward specific types of responses, can potentially reveal sensitive information, e.g., their preferences for some specific political parties. Therefore, we augment our RL-LOW with a label-DP protection mechanism to mitigate the risk of privacy breaches from the labels.
A.2 A Motivational Example of MDP with Known Transition Probabilities
The MDP setting of our Section 5, in which transition probabilities are assumed to be known, is inspired by the inference framework of most language models. In this framework, states correspond to prompts, and actions represent the subsequent response or token. For instance, consider the current state: "The capital of France is". If the action "Paris" is taken, the subsequent state is deterministically defined as "The capital of France is Paris". While the reward associated with this transition may be unknown, the transitions themselves are deterministic and well-defined in this context. Therefore, the assumption of knowing transition probabilities aligns with the behaviour of language models, where the next state is known based on the current state and action, and this motivates the MDP setting of our Section 5.
Appendix B Numerical Simulations
@@
In this section, we present our numerical simulations of our algorithm RL-LOW and and its differentially private counterpart DP-RL-LOW. We compare them to the state-of-the-art (non-private algorithm) Pessimistic MLE developed by Zhu et al. (2023). We conduct the experiments on a synthetic dataset. Specifically, we set the number of states $S=2$, the number of actions $A=10$, the dimensionality of the data $d=5$, the unknown parameter vector $\theta=[1,1,1,1,1]^{\top}$, and the state distribution $\rho=[0.4,0.6]$. The feature vector of each action is generated as follows: For the $i$-th action of state $k\in\{1,2\}$, we first uniformly generate a $d$-dimentional vector $\phi^{\prime}(k,i)$ with all non-negative elements and $\lVert\phi^{\prime}(k,i)\rVert_{1}=1$. Then, for each state $k\in\{1,2\}$, we set the feature vector of $i$-th action as $\phi(k,i)=\phi^{\prime}(k,i)-0.01(i-1)\theta$. That is, in both state $1$ and $2$, the best action is the first action, and the suboptimality gap of the $i$-th action is $0.05i$. In addition, for both states $k\in\{1,2\}$ and $i<j$, we set $N_{k,i,j}=\frac{1}{A(A-1)}$, i.e., the proportions of comparisons for this instance are uniform. In the simulation, we use $\lceil N_{k,i,j}n\rceil$ as the number of samples in the comparison between actions $i$ and $j$ under state $k$.
@@
As for the hyperparameters of Pessimistic MLE, we follow the default setting of Zhu et al. (2023, Section 3). In addition, Pessimistic MLE only works under the assumption that $\langle 1,\theta\rangle=0$ (Zhu et al. 2023, Assumption 2.1). Therefore, when running the experiments of Pessimistic MLE, we further set $d=6$, $\theta=[1,1,1,1,1,-5]$ and the $6$-th element of each feature vector is set to 0. Then, this new instance is mathematically equivalent to the original instance and additionally satisfies Assumption 2.1 of Zhu et al. (2023) which is needed for Pessimistic MLE.
@@
Lastly, as shown in Figure 1 (right) and the curve of DP-RL-LOW of Figure 1 (left), it is evident that achieving comparable performance between RL-LOW and DP-RL-LOW may necessitate substantially larger sample sizes $n$ when considering small privacy parameters of $\varepsilon$ and $\delta$. This observation is again consistent with our theoretical findings in Theorem 6.3.
Figure 1: Comparison of RL-LOW and DP-RL-LOW to Pessimistic MLE on average simple regret and standard deviation (shaded area). In the left figure, we set $\delta=0.2$ and $\varepsilon=0.9$ for DP-RL-LOW. In the right figure, we set $n=400$ for all policies.
Appendix C More Explanations for the Hardness Parameter $H(v)$
The hardness parameter $H(v)$ is inversely proportional to the square of the suboptimality gaps across various states and actions. Specifically, a smaller suboptimality gap implies an increased hardness parameter, indicating that the instance is more challenging to learn. This relationship underscores the significance of the suboptimality gap as a critical measure in evaluating the complexity and learning difficulty of each instance.
For example, let’s suppose $S=1,A=2$, $r_{1,1}=0$ and $r_{1,2}=1$. Given offline data with of size $n$ (i.e., these are $n$ history records for "action $1$ wins action 22" or "action $2$ wins action 11"), if the learner picks the action with the most winning records (as it does in our algorithm RL-LOW), then by Hoeffding’s inequality, it will suffer an upper bound of expected simple regret of $\exp(-C\cdot\frac{n}{{(r_{1,2}-r_{1,1})^{-2}}})$ for a constant $C$ that does not depend on $r_{1,2}-r_{1,1}$ (in fact this upper bound is also tight in the hardness parameter according to our lower bound). Notice the exponent is $\Theta(\frac{-n}{{(r_{1,2}-r_{1,1})^{-2}}})$, and the hardness parameter $H(v)$ is exactly $(r_{1,2}-r_{1,1})^{-2}$ under this instance.
Appendix D Useful facts
@@
Definition D.1 (SubGaussian norm).
@@
Definition D.2 (Variance proxy).
@@
Lemma D.3 (Linear combination of subGaussian random variables).
Let $X_{1},\ldots,X_{n}$ be independent subGaussian random variables, where the variance proxy of $X_{i}$ is $\sigma_{i}^{2}$ . Then, for any $a_{1},\dots,a_{n}\in\mathbb{R}$, the random variable $\sum_{i=1}^{n}a_{i}X_{i}$ is a subGaussian random variable with variance proxy $\sigma^{2}=\sum_{i=1}^{n}a_{i}^{2}\sigma_{i}^{2}$.
Lemma D.4.
@@
Lemma D.5.
(Adapted from Vershynin (2018, Proposition 2.5.2)) For any subGaussian random variable $X$,
@@
Appendix E Proof of Proposition 2.3
Lemma E.1.
@@
which is a contradiction to that $v$ is inconsistent. This completes the proof of Lemma E.1. ∎
@@
Fix any inconsistent instance $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$. By Lemma E.1, there exists $(\underline{k},\underline{i})\in\mathcal{S}\times\mathcal{A}$ with $\underline{i}\neq i^{*}_{\underline{k}}$ such that
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{v}^{\pi}\left[\sum_{k\in\mathcal{S}}\rho_{k}\left(\max_{j\in\mathcal{A}}\langle\phi(k,j)-\phi(k,\hat{i}_{k}),\theta\rangle+\max_{j\in\mathcal{A}}\langle\phi(k,j)-\phi(k,\hat{i}_{k}),\theta^{\prime}\rangle\right)\right]$ (27)
@@
Appendix F Proof of Proposition 3.2 and More Details on Computational Complexity
F.1 Proof of Proposition 3.2
@@
Lemma F.1.
Fix any consistent instance $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$. Recall the definitions of $w^{(k,i,j)}$ and $\gamma_{k,i,i^{*}_{k}}$ in Eqn. (7) and Eqn. (3.1), respectively. Then, for any $(k,i)\in\mathcal{S}\times\mathcal{A}$ with $i\neq i^{*}_{k}$,
@@
$\phi(k,i)-\phi(k,j)]_{\mathcal{G}}=\sum_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}u_{k^{\prime},i^{\prime},j^{\prime}}[\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime})]_{\mathcal{G}}.$
@@
By Lemma F.1, we get that under RL-LOW, for any $(k,i,j,j_{2})\in\mathcal{S}\times\mathcal{A}^{3}$,
@@
which implies that $\lvert\hat{\mathcal{I}}_{k}\rvert\geq 1$ and
$\operatorname*{argmax}_{i\in\mathcal{A}}\hat{r}_{k,i,j_{1}}=\operatorname*{argmax}_{i\in\mathcal{A}}\hat{r}_{k,i,j_{2}}=\hat{\mathcal{I}}_{k}\quad\mbox{for any}\quad j_{1},j_{2}\in\mathcal{A}.$ (34)
This completes the proof of Proposition 3.2. ∎